A door delivery florist wishes to estimate the proportion of people in his city that will purchase his flowers. Suppose the true proportion is 0.05 . If 300 are sampled, what is the probability that the sample proportion will differ from the population proportion by more than 0.03 ? Round your answer to four decimal places.
Answer #1
for normal distribution z score =(p̂-p)/σp
here population proportion=     p= 0.0500
sample size       =n= 300
std error of proportion=σp=√(p*(1-p)/n)= 0.0126
probability =1-P(0.02<X<0.08)=1-P((0.02-0.05)/0.013)<Z<(0.08-0.05)/0.013)=1-P(-2.38<Z<2.38)=1-(0.9913-0.0087)=0.0174